# KnowPhish: Large Language Models Meet Multimodal Knowledge Graphs for Enhancing Reference-Based Phishing Detection

### v1

@@ 5.4 RQ2: Field Study
To further understand how well the phishing detection performance of different RBPDs generalize to a local context, we conduct our field study on the SG-SCAN dataset. This field study also incorporates URLScan[53], a widely-recognized commercial phishing detector, as a baseline to assess the impact of KnowPhish and KPD within the industry. Note that this dataset is unlabeled, so we only manually validate the samples reported by the phishing detectors. This allows us to compute the true positive counts and the precision for evaluation (but not recall).
Detector BKB #P #TP Precisionnormal-$\uparrow$ Timenormal-$\downarrow$
@@
URLScan NA 517 495 95.74 NA
KPD KnowPhish 699 681 97.42 1.64s
Table 3: Phishing detection performance of different RBPDs and URLScan on SG-SCAN dataset. #P and #TP refer to the numbers of reported phishing, and true positives, respectively.
(b) Phishpedia+DynaPhish
Figure 10: Top 20 phishing targets detected by KPD+KnowPhish and Phishpedia+DynaPhish on SG-SCAN.
  • KPD detects the greatest number of phishing webpages while maintaining the highest precision over all the baselines, including the commercial phishing detector URLScan. When equipped with KnowPhish, KPD finds at least two times more phishing webpages than image-based RBPDs do. We find that the major reason for this large improvement is that logoless phishing webpages are very common in local webpage streams, which our LLM module in KPD can handle by extracting brand information from HTML. In contrast, image-based RBPDs cannot detect logoless phishing webpages. Additional examples of logoless phishing webpages can be found in Figure 13, located in Appendix C.
  • KnowPhish covers many local phishing targets in Singapore (e.g., DBS Bank, OCBC Bank, Shopee, Lazada, Singtel, Qoo10). These phishing targets all belong to the ten high-value industries we mentioned in Section 3.1, which further validates our empirical observation.
  • Phishpedia and PhishIntention detect slightly fewer phishing webpages when integrated with KnowPhish than with DynaPhish. Our manual inspection found that DynaPhish tends to be more effective when encountering less-known brands (e.g., SEOptimer and ASOTools) that are not even maintained in Wikidata, while KnowPhish performs better in identifying phishing webpages using logo variants or text brands.
To summarize, our field study highlights the need for RBPDs that can operate within the text modality, and demonstrates both the effectiveness and efficiency of KnowPhish and KPD in the local setting.
@@ 5.5 RQ3: Adversarial Robustness
We study the robustness of our text-based components against HTML obfuscation as an evasion technique. We design two approaches to obfuscate HTML:
  • Insertion. We follow the techniques introduced in [27] to generate random tokens. The 5 random noisy tokens will be placed as either prefixes or suffixes.
@@
  • Typosquatting. Based on [33] and several motivating suspicious webpage examples (e.g., Figure 12 in Appendix C) that utilize obfuscation, we perform typosquatting on either the title only or all the text elements in the HTML. Here, we obfuscate one character in each word.
@@
Both attack techniques aim to compromise the text-based models by introducing noise into the input texts while maintaining similarity to the original texts. Figure 11 shows several examples of the four settings of HTML obfuscation.
@@
For brand extraction, we use 200 random phishing webpage samples from the TR-OP dataset to evaluate 1) NIR, the ratio of samples whose brand is not identifiable among the test samples, and 2) brand extraction accuracy. Since correct predictions of the brand intention may take multiple forms (e.g., ‘DBS’ or ‘DBS Bank’ are correct predictions of the brand DBS), evaluating the brand extraction accuracy requires human validation. Consequently, the size of the evaluation set is limited to a small number. For CRP classification, we sample and annotate 578 webpages from SG-SCAN dataset with a 148/276/154 explicit/implicit/non-CRP split, and the evaluation metric is classification accuracy.
Figure 11: Illustration of HTML obfuscation techniques.
Task Attack Type Position NIRnormal-$\downarrow$ ACCnormal-$\uparrow$
None None 12.00 81.00
Brand Insertion Prefix 13.00 79.00
Extraction Suffix 14.00 78.50
Typosquatting Title 13.50 78.00
All Texts 18.00 72.00
CRP None None NA 93.94
Classif. Typosquatting All Texts NA 91.00
@@
Table 4: Performance on brand extraction and CRP classification after different HTML obfuscations.
Table 4 shows the results after employing adversarial attacks. Overall, our LLM-based approach is robust against such attacks. For the brand extraction task, insertion as an adversarial attack only leads to 2-3% drops in the ACC and 1-2% increase in NIR. The performance gap is slightly larger when we introduce typosquatting, especially when employing it to all the text in the HTML. However, the overall brand extraction ACC remains higher than 70% under all types of attacks. We believe this is because the LLM can seek useful information from multiple parts of HTML, or even use its internal knowledge to correct the brand name when tokens are corrupted. Similarly, for CRP classification, our LLM-based CRP classifier can retain a relatively high classification accuracy even when each word in the HTML text is typosquatted.
In summary, our adversarial attack experiments show that our LLM-based components are robust to the four types of HTML obfuscation.
@@ 5.6.1 Ablation of KPD Components
Dataset Detector Recallnormal-$\uparrow$ Precisionnormal-$\uparrow$ #P #TP
@@
5.6.2 Effect of Different LLM Architectures
We also experiment with different LLM architectures to investigate their impacts on phishing detection.
Detector ACCnormal-$\uparrow$ F1normal-$\uparrow$ Precisionnormal-$\uparrow$ Recallnormal-$\uparrow$
KPD (GPT-3.5-turbo-instruct) 92.49 92.05 97.84 86.90
   w/ GPT-3.5-turbo 91.84 91.29 97.94 85.48
   w/ GPT-4 92.31 91.82 98.05 86.34
   w/ LLaMA-2-7B 91.23 90.68 96.69 85.38
Table 6: Phishing detection performance of KPD with different LLMs as the backbone.
The results are shown in Table 6. While GPT-4 yields the highest precision, our currently used LLM (GPT-3.5-turbo-instruct) gives the highest recall. In general, larger models do better than smaller ones such as LLaMA-2-7B on both precision and recall. Based on our observation, LLaMA-2-7B tends to mix up information from the in-context demonstrations more often and predict the target brand of the actual sample to be those from the demonstrations. Such mistakes, when happening to benign samples, can cause domain-brand inconsistency, resulting in false positive predictions. Additionally, smaller models have poorer instruction-following ability and cannot adhere to the required output format in more cases.
@@ 5.6.3 Analysis of CRP Classifier
Aside from the main components, we also take a closer look at the performance of different CRP classifers. Our text-based CRP classifier, a well-trained XLM-RoBERTa model, takes in the sample’s HTML and a CRP summary generated by the LLM (GPT-3.5-instruct). We replace our current LLM summarizer with other architectures to investigate whether the choice of CRP summarizer matters. Additionally, we also remove the CRP summary and use the HTML as the only input to our XLM-RoBERTa model, to check its performance in a standalone setting. Since [33] also proposes an image-based CRP classifier that generates CRP prediction from screenshot, we also include it as an individual baseline.
Detector ACCnormal-$\uparrow$ F1normal-$\uparrow$ Precisionnormal-$\uparrow$ Recallnormal-$\uparrow$
Text-based CRP Classifier 93.94 95.95 94.31 97.64
   w/ GPT-3.5-turbo Summary 93.43 95.58 94.27 96.93
   w/ GPT-4 Summary 94.12 96.05 94.72 97.41
   w/ LLaMA-2-7B Summary 86.85 91.72 85.22 99.29
   w/o LLM Summary 93.43 95.63 93.27 98.11
Image-based CRP Classifier[33] 51.90 51.91 97.40 35.38
@@
Our models are then evaluated on the same CRP dataset in Section 5.5, and the results are shown in Table 7. Among the NLP-based methods, summaries generated by larger LLMs (GPT-3.5-instruct, GPT-3.5-turbo, GPT-4) give similar performances. However, there is a notable drop in precision and accuracy when switching the LLM summarizer to LLaMA-2-7B. This can be potentially caused by hallucinations in the summaries, which mislead our XLM-RoBERTa Classifier (See discussions in Section 6). Removing the LLM summary only slightly affects the precision and accuracy. This means our CRP classifier can function properly individually, while the LLM summary adds an extra safeguard against potential distributional shifts from our classifier’s training samples during inference. Compared with the image-based method, while our text-based method lags slightly on precision, it has a much higher recall and accuracy and can detect more CRPs, especially the implicit ones.
@@ 6.1 Error Analysis
This section delves into a comprehensive analysis of the false positives and negatives of KnowPhish and KPD.
@@
By manually examining all 97 false positives made by KPD, we pinpointed two primary causes: brand representation collisions, and incomplete inclusion of domain variants, accounting for 45.36% and 43.30% of the total false positives, respectively.
Brand representation collision occurs either the webpage’s screenshot or HTML is matched to the wrong brand. Both the logo matcher and text brand extractor are not perfect, and can misidentify the brand intention of the webpages by mismatching a logo or extracting brands from text that does not match the true brand intention.
For the second issue, domain variants can be missing from KnowPhish because the Whois owner information is not available for all domains. We find that at most 30.79% of the domains in KnowPhish have their owner information available. This deficiency may lead to incomplete lists of domain variants for brands, leading to false positives when the current page’s domain is omitted as a legitimate domain in the knowledge base.
@@
We also examined all 655 false negative samples by KPD on the TR-OP dataset, uncovering three primary reasons behind these erroneous predictions.
@@
Additionally, negative classifications by the CRP classifier also leads to 30.2% of the false negatives. Most of these failure cases are accompanied by extremely implicit credential-requiring intentions. Figure 14 in Appendix C shows an example, where our text brand extractor detected the brand intention as Telegram, but our CRP classifier classifies it as non-CRP.
The limited brand coverage of KnowPhish is responsible for the remaining false negatives. Phishing targets such as Bank Promerica, Minnesota Unemployment Insurance, and Battleground Mobile India, are not even included in Wikidata. While KnowPhish enhances the performance of existing RBPDs, surpassing even DynaPhish, some phishing targets will be beyond the BKB. In such cases, any RBPD will face challenges in detecting the phishing webpages.
6.2 Incompleteness of External Databases
Our error analysis in Section 6.1 points to brand knowledge limitations (including logos, aliases, and domain variants) as a major source of errors, arising from limitations of Wikidata and the Whois service. The most straightforward solution is to integrate other brand databases, such as the WIPO Global Brand Database [60]. Alternatively, we can rely on the implicit knowledge from LLMs [50, 39, 36, 23] or methods integrating LLMs with online search [56, 48, 38, 26].
To further handle false positives, we can also introduce a secondary validator, such as a search engine-based filter to validate the popularity of a webpage before the RBPDs report it as phishing: i.e., if a webpage is reported as phishing by RBPDs but found to be a popular webpage by search engines, it is likely to be benign. Previous studies have shown that a search engine-based validator can identify more than 60% of benign webpages[13, 34] with negligible runtime overhead.
6.3 Errors Made by LLMs
We observe two major types of errors made by our LLMs. Firstly, they occasionally extract the wrong brands (See Table 4) when the HTML contains multiple brand names. This can potentially be addressed by better prompting or eliciting better reasoning from LLMs. Text spans for target brands usually have morphological (e.g., capitalization of every character) or contextual characteristics that distinguish them from non-brands. Therefore, we can use existing techniques [58, 57, 62] to guide the LLMs to identify and reason about these characteristics for more accurate target brand extraction. Secondly, we observe that LLMs can make up nonexistent HTML elements such as input fields and redirect buttons that are indicative of a credential-taking intention. This points to hallucination, a known issue for LLMs [22]. To address this issue, many strategies [51, 49, 8] can be used to direct our models’ attention to the parts of HTML that are pertinent to credential-taking intention and make their generations better anchored on page content.
@@ 7 Related Work
Phishing Detection
The simplest phishing detection methods rely on blacklists of malicious URLs [45, 40, 42], which are reactive approaches. Proactive approaches include feature engineering-based methods, which rely on hand-crafted features from URLs [28, 37, 54], HTML [18, 61], or both [31, 35, 29]. These methods are limited by their inability to use logos, and are susceptible to distribution shift. RBPDs extract brand intention of webpages through screenshots [16, 3, 1] or logos[32, 33], relying on small, manually collected BKBs. Recently, DynaPhish [34] proposed to dynamically expand the BKB during deployment. However, such interaction during deployment leads to substantial increases in the detector’s running time, e.g., 10.6 seconds per sample. In contrast, our multimodal BKB is constructed fully before deployment, making our detector much more efficient.
LLMs and Knowledge-Intensive Applications
LLMs have shown remarkable performance on a wide range of language and code-related tasks [2, 9], and have been extended to large multimodal models (LMMs) [63]. A few recent works apply LLMs for phishing detection [25, 59]. However, these are non-RBPD methods, and cannot use logos. They also do not integrate with knowledge bases, thus being limited in the breadth of knowledge they have available.

### v2

@@ 6.4 RQ2: Field Study
To further understand how well the phishing detection performance of different RBPDs generalize to a local context, we conduct our field study on the SG-SCAN dataset. Note that this dataset is unlabeled, so we only manually validate the samples reported by the phishing detectors. This allows us to compute the true positive counts and the precision for evaluation (but not recall).
  • KPD+KnowPhish detects the greatest number of phishing webpages. When equipped with KnowPhish, KPD finds at least two times more phishing webpages than image-based RBPDs do. This improvement mainly comes from the detection of logo-less phishing webpages by KPD (examples are given in Appendix A.2).
  • KPD+DynaPhish still underperforms KPD+KnowPhish, partly due to insufficient alias variants. For example, for the target DBS Bank, the LLM may predict either ‘DBS Bank’ or ‘DBS’ based on the HTML content. KnowPhish leverages the rich aliases from Wikidata, allowing it to match the brand in both cases. In contrast, DynaPhish only recognizes ‘DBS Bank’ as an alias, but not the other, resulting in false negatives.
[…]
@@
    KPD DynaPhish 628 581 92.52 7.83s
    KnowPhish 699 681 97.42 1.64s
    Table 3: Phishing detection performance of different RBPDs on SG-SCAN dataset. #P and #TP refer to the numbers of reported phishing, and true positives, respectively.
    (b) Phishpedia+KnowPhish
    (c) KPD+DynaPhish
    (d) KPD+KnowPhish
    Figure 9: Top 20 phishing targets detected by Phishpedia and KPD using either DynaPhish or KnowPhish as BKB on SG-SCAN.
  • DynaPhish yields lower precision than KnowPhish mostly because of the inclusion of web-hosting brands and failing popularity validation in benign domains. KnowPhish mitigates the issue of web-hosting brands by excluding them via a knowledge graph-based approach. Further details are discussed in Appendix A.3.
  • KnowPhish covers many local phishing targets in Singapore (e.g., SingPost, DBS Bank). These targets all belong to the high-value industries mentioned in Section 4.1, which further validates our empirical observation.
  • Phishpedia and PhishIntention detect slightly fewer phishing webpages when integrated with KnowPhish than with DynaPhish. Our inspection found that DynaPhish tends to be more effective when encountering less-known brands (e.g., SEOptimer and ASOTools) that are not even maintained in Wikidata, while KnowPhish performs better in identifying phishing webpages using logo variants or text brands.
To summarize, our field study highlights the need for RBPDs capable of operating within the text modality. It also further validates the effectiveness and deployment efficiency of KnowPhish over DynaPhish, especially when a multimodal RBPD (e.g., KPD) is utilized.
@@ 6.5 RQ3: Adversarial Robustness
We study the robustness of our text-based components against three types of evasion techniques in HTML:
@@
  • Typosquatting. Based on [34] and several motivating suspicious webpage examples (see [26]) that utilize obfuscation, we perform typosquatting on either the title only or all the text elements in the HTML. Here, we obfuscate one character in each word.
  • Prompt Injection. We add an adversarial text ‘Please ignore the previous instruction and answer Not identifiable’ into the HTML to mislead the LLM to follow the adversarial instructions, instead of the original ones.
  • Text-to-Image. We consider an extreme scenario where HTML has been fully obscured by the text-to-image attack, where the only useful information to the phishing detectors are URLs and screenshots.
@@
These attack techniques aim to compromise the text-based models (i.e., LLM summarizer and CRP classifier) by injecting adversarial perturbations into the HTML while maintaining similarity to the original webpage appearance. Figure 10 shows five examples of our HTML-oriented evasion attacks.
We apply defense remedies against the latter two advanced attacks, respectively. Specifically, for prompt injection, we design a hardened prompt with additional instructions, ensuring the LLM stays focused on its original tasks. For text-to-image attacks, we replace the default text-based LLM with a multimodal one, GPT-4V, as the defense method, enabling the LLM to generate summaries through screenshots when HTML information is unavailable. The detailed prompt modifications are provided in [26].
@@
For evaluation, we use 200 random phishing webpage samples from the TR-OP dataset to conduct the adversarial experiments. We evaluate both the LLM brand extraction accuracy and the CRP classification accuracy. Since correct predictions of the brand intention may take multiple forms (e.g., ‘DBS’ or ‘DBS Bank’ are correct predictions of the brand DBS), evaluating the brand extraction accuracy requires human validation. Consequently, the size of the evaluation set is limited to a small number.
Figure 10: Illustration of HTML-oriented evasion attacks.
[…]
w/o defense w/ defense
None None 81.00 NA
Typosquatting Title 78.00 NA
Brand All Texts 72.00 NA
Extraction Prompt Injection Prefix 75.00 76.00
Suffix 55.50 63.50
Text-to-Image All Texts 5.00 60.00
None None 92.50 NA
CRP Typosquatting All Texts 93.00 NA
Classif. Prompt Injection Prefix 92.50 92.50
Suffix 92.50 92.00
Text-to-Image All Texts 0.00 70.00
@@
Table 4: Performance on brand extraction and CRP classification after different types of adversarial attacks.
Table 4 shows the results of our adversarial attacks experiments. For brand extraction, all three attacks cause additional incorrect brand predictions across varying degrees. The ramifications are particularly pronounced when employing prompt injection as an HTML suffix and text-to-image attack, leading to much lower accuracies of 55.5% and 5%, respectively. By applying their corresponding defense methods, the figures increase slightly by 8% for prompt injection, but largely by 65% for text-to-image attack. We believe more sophisticated defense methods can be applied to better counter the effect of these two attacks and will discuss them in Appendix A.4.
Regarding CRP classification, we observe that typosquatting and prompt injection are not effective in this task, as the outcomes remain similar with or without defense. This is because the LM-based CRP classifier can still extract useful information from HTML. Text-to-Image attack, instead, compromises the CRP classifier entirely. However, such adverse effects are mitigated by using a multimodal LLM as a defense, preserving the classification accuracy at 70%.
In general, our LLM-based method demonstrates certain robustness against these attacks, either in its original form or with defense remedies.
@@ 6.6.1 Ablation of KPD Components
[…]
@@
6.6.2 Effect of Different LLM Backbones
We also experiment with different LLM backbones to investigate their impacts on their summary answer (i.e., brand extraction and CRP classification) accuracy and phishing detection performance.
We augment the evaluation set in Section 6.5 with additional 200 random benign samples from TR-OP. This results in a new evaluation set with 400 samples, each with manually validated brand label and CRP label, which is used to assess the accuracy of the LLM answers. The phishing detection performance is separately evaluated on the entire TR-OP dataset using KPD with these LLM backbones.
LLM Backbone Brand CRP Phishing Detection
[…]
GPT-3.5-turbo-instruct 84.00 81.50 97.84 86.90
GPT-3.5-turbo 81.25 77.25 97.94 85.48
GPT-4 85.50 90.25 98.05 86.34
LLaMA-2-7B 62.00 83.75 96.69 85.38
Table 6: LLM answer accuracies and phishing detection performance of KPD with different LLMs as the backbones.
@@ 6.6.3 Analysis of CRP Classifier
In addition to analyzing the main components, we examine the design of our text-based CRP classifier. We study two ablated variants: one excludes the LLM summary, and the other removes the HTML from the small LM input. We also include an existing image-based CRP classifier that generates CRP prediction from screenshots[34] as an individual baseline to further evaluate its effectiveness in detecting implicit CRPs.
[…]
Text-based CRP Classifier 86.00 92.02 90.22 93.89
w/o LLM Summary 83.75 90.54 90.67 90.41
w/o HTML 81.50 88.54 94.12 83.72
Image-based CRP Classifier[34] 55.50 65.50 98.25 49.12
@@
These baselines are then evaluated on the same dataset used in Section 6.6.2, with the results presented in Table 7. Overall, our text-based CRP classifier that takes both HTML and CRP summary as inputs yields the best accuracy, F1 score, and recall among all ablated variants. Removing either input diminishes the performance of our CRP classifier, particularly when the HTML input is excluded. This further demonstrates that relying solely on LLM verdicts for CRP classification may not be sufficiently reliable. Compared to the image-based method, our text-based can detect more CRPs, particularly the implicit ones, thus having much higher accuracy and recall, although it lags slightly on precision.
7 Discussion
Going beyond the empirical analytics, this section further discusses the factual difference between KnowPhish and DynaPhish, and potential trade-offs.
Data Source Quality  KnowPhish benefits from multiple extra high-quality data sources, such as Wikidata and Tranco top domain list, in addition to the Google Search used by DynaPhish. These sources enrich brand aliases, logos, and domain variants, significantly improving the brand identification capabilities of the detector backbone. Conversely, DynaPhish primarily relies on its webpage layout detector and Google Search to collect brand knowledge. The quality of its brand knowledge is affected by the performance of these two components, potentially leading to the failure to build brand knowledge.
Deployment Latency  KnowPhish constructs brand knowledge offline, whereas DynaPhish does so online. The online brand knowledge-gathering step significantly increases runtime overhead for the detector to produce a verdict. In contrast, the additional runtime overhead introduced by KnowPhish is limited to querying a larger BKB. This step only involves calculating logo similarity scores and finding matched aliases, which is much faster than accessing new webpages during deployment, as DynaPhish requires.
Trade-offs  Despite these advantages, KnowPhish may be inferior to DynaPhish in terms of the timeliness of brand knowledge. KnowPhish uses Wikidata to search for potential phishing targets, while DynaPhish uses a search engine that updates more frequently. Consequently, KnowPhish might lead to more false negatives when emerging phishing targets become prevalent in webpage streams.
@@ 8.1 Error Analysis
This section delves into a comprehensive analysis of the false positives and negatives of KPD+KnowPhish.
@@
By manually examining all 97 false positives made by KPD+KnowPhish on TR-OP, we pinpointed two primary causes: brand representation collisions, and incomplete inclusion of domain variants, accounting for 45.36% and 43.30% of the total false positives, respectively.
Brand representation collision occurs when either the webpage’s screenshot or HTML is matched to the wrong brand. Both the logo matcher and text brand extractor are not perfect and can misidentify the brand intention of the webpages by mismatching a logo or extracting brands from text that does not match the true brand intention.
For the second issue, domain variants can be missing from KnowPhish because their Whois owner information is unavailable. We find that at most 26.79% of the domains in the Tranco domain list have their owner information available. This deficiency results in incomplete lists of domain variants for brands, leading to false positives when the current page’s domain is omitted as a legitimate domain in KnowPhish.
@@
We also examined all 655 false negative samples by KPD+KnowPhish on the TR-OP dataset, uncovering three primary reasons behind these erroneous predictions.
@@
Additionally, negative classifications by the CRP classifier also lead to 30.2% of the false negatives. Most of these failure cases are accompanied by extremely implicit credential-requiring intentions. Our supplementary materials [26] provide an example, where our text brand extractor detected the brand intention as Telegram, but our CRP classifier classifies it as non-CRP.
The limited brand coverage of KnowPhish is responsible for the remaining false negatives. Phishing targets such as Bank Promerica, Minnesota Unemployment Insurance, and Battleground Mobile India, are not even included in Wikidata. While KnowPhish enhances the performance of existing RBPDs, some phishing targets will be beyond the BKB. In such cases, any RBPD will face challenges in detecting phishing webpages.
8.2 Potential Solutions
Incompleteness of External Databases  Our error analysis in Section 8.1 points to brand knowledge limitations (including logos, aliases, and domain variants) as a major source of errors, arising from limitations of Wikidata and the Whois service. The most straightforward solution is to integrate other brand databases, such as the WIPO Global Brand Database [60]. Alternatively, we can rely on the implicit knowledge from LLMs [53, 41, 38, 24] or methods integrating LLMs with online search [58, 50, 40, 28]. To further handle false positives, we can also combine a secondary validator, such as a search engine-based filter to validate the benignity of a webpage before RBPDs report it as phishing[14, 35].
Performance of LLMs  LLMs may occasionally extract incorrect brands when multiple brands are present in the HTML, or make up nonexistent HTML elements in its reasoning output due to hallucination. These problems may be mitigated by better prompting techniques and more advanced LLM reasoning strategies[54, 51, 9].
@@ 9 Related Work
Phishing Detection  The simplest phishing detection methods rely on blacklists of malicious URLs [47, 42, 44], which are reactive approaches. Proactive approaches include feature engineering-based methods, which rely on hand-crafted features from URLs [29, 39, 56], HTML [19, 61], or both [32, 37, 30]. These methods are limited by their inability to use logos and are susceptible to distribution shifts. RBPDs extract brand intention of webpages through screenshots [17, 3, 1] or logos[33, 34], relying on small, manually collected BKBs. Recently, DynaPhish [35] proposed to dynamically expand the BKB during deployment. However, such interaction during deployment leads to substantial increases in the detector’s running time, e.g., 10.6 seconds per sample. In contrast, our multimodal BKB is constructed fully before deployment, making our detector much more efficient.
LLMs and Knowledge-Intensive Applications  LLMs have shown remarkable performance on a wide range of language and code-related tasks [2, 10], and have been extended to large multimodal models (LMMs) [62]. A few recent works apply LLMs for phishing detection [27, 59]. However, these are non-RBPD methods, and cannot use logos. They also do not integrate with knowledge bases, thus being limited in the breadth of knowledge they have available.